OpenStudy (anonymous):

Help me asap !! i.imgur.com/TBBIaib.jpg i.imgur.com/SJJmL7B.jpg Someone help me with the questions above?? Make answers simple please.

OpenStudy (britanyashton):

alrighty! I'll help ya! =]

OpenStudy (britanyashton):

Well, do you remember the formula y=mx+b? What does b represent in the formula, and which letter represents the slope in the formula I just wrote?

OpenStudy (britanyashton):

Write that down here, and I'll tell ya what to do from there. Is graphing a part that you understand, or would you like help with that too? =]

OpenStudy (britanyashton):

The answer to the solution of equations = You'll get this, when you graph both of those. The place where they intersect, will be the solution(s).

OpenStudy (anonymous):

I need help with that, too. B is the Y-intercept and M is the slope of the line ??

OpenStudy (britanyashton):

Yes!! =] Rise/run is slope, when graphing it. We have a negative slope here.

OpenStudy (britanyashton):

The slope for graphing is -1/2....sooo down one, right 2. =)

OpenStudy (britanyashton):

start at (0,0) and do that on your graph. You'll get a straight line.

OpenStudy (britanyashton):

Does that make sense?

OpenStudy (britanyashton):

Here's some graphing paper.

OpenStudy (anonymous):

Uhh.. you want me to make a straight line from (0,0?)

OpenStudy (britanyashton):

Here, lemme explain it this way. y=mx+b y=-1/2x+0 (i'm adding in that zero, because it's understood when none is written) b is the y intercept, like where it intersects the y (or vertical) line/access. m is the slope. there is a language for slope, in some words, which is RISE/RUN. Rise is up and down, run is how much left or right. In this problem you have here, the slope is -1/2 the "rise" of the slope m, is negative, so therefore it is down 1. the "run" is two here, a positive two, so its up two. We know a point here - the y-intercept! The slope is basically, in what direct, and how steep it is. rise/run is used when you're starting at your y intercept, in this case (0,0) then going -1/2 or down 1 from that point, and then right two. keep doing that each point. lemme show u.

OpenStudy (britanyashton):